# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 5.2.2 Frequency Sampling Method
The Frequency Sampling Method designs FIR filters by specifying the desired frequency response at a set of equally spaced frequency points and then computing the corresponding time-domain coefficients using the inverse DFT [155].
@@
    Select the Number of Frequency Samples: Choose $N$, the length of the filter, which determines the number of frequency samples [136].
@@
    Specify the Desired Frequency Response: Define the desired frequency response $H_{d}[k]$ at the $N$ frequency points [136].
@@
    Compute the Time-Domain Coefficients: Obtain the filter coefficients $h[n]$ by computing the inverse DFT of $H_{d}[k]$ [155]:
@@
    Implement the Filter: Use the calculated coefficients $h[n]$ in the FIR filter implementation [155].
@@ 5.4 IIR Filter Design Techniques
Infinite Impulse Response (IIR) filters are widely used in digital signal processing because of their efficiency and ability to provide sharp frequency responses with a relatively low filter order [184]. Unlike FIR filters, IIR filters have feedback, which gives them an infinite impulse response. In this section, we will discuss two fundamental techniques for designing IIR filters: the bilinear transform method and the impulse invariant method [136].
@@ 5.4.1 Bilinear Transform Method
The bilinear transform is one of the most popular methods for designing IIR filters. It allows us to take an analog filter design (such as a Butterworth, Chebyshev, or Elliptic filter) and map it into the digital domain [155]. This method works by mapping the analog $s$-plane to the digital $z$-plane using the following transformation:
@@
The bilinear transform preserves the frequency response characteristics of the analog filter while avoiding aliasing (a problem where high frequencies in the analog system would otherwise appear as lower frequencies in the digital system) [184]. However, one drawback is frequency warping, where the relationship between analog and digital frequencies becomes nonlinear. This can be corrected by pre-warping the critical frequencies before applying the transformation [136].
@@
18 plt.xlabel(’Normalized Frequency (×π rad/sample)’)
@@ 5.4.2 Impulse Invariant Method
The impulse invariant method is another technique for converting analog filters into digital filters [155]. The key idea behind this method is to match the impulse response of the analog filter to the impulse response of the digital filter. This technique is particularly effective for designing low-pass filters where it is important to retain the shape of the analog impulse response in the digital domain [184].
In the impulse invariant method, we first compute the impulse response of the analog filter by taking the inverse Laplace transform of its transfer function. We then sample this continuous-time impulse response to obtain the discrete-time impulse response, which is subsequently used to determine the digital filter’s transfer function [136].
One major drawback of the impulse invariant method is that it can introduce aliasing because high-frequency components of the analog filter may be folded into lower frequencies in the digital domain [184]. Therefore, this method is best suited for low-pass filters where aliasing is less of a concern [155].
@@
17 plt.xlabel(’Normalized Frequency (×π rad/sample)’)
@@ 5.4.3 Comparison of Bilinear Transform and Impulse Invariant Methods
Both the bilinear transform and impulse invariant methods have their advantages and drawbacks. Understanding their differences will help determine when to use each method [136].
  • Bilinear Transform: The bilinear transform avoids aliasing by warping the frequency axis. However, it introduces frequency warping, which must be compensated for, especially in high-frequency regions [155].
  • Impulse Invariant Method: The impulse invariant method preserves the shape of the analog filter’s impulse response, but it can introduce aliasing. This method is typically preferred for low-pass filters where aliasing is less of a concern [184].
@@
In most practical applications, the bilinear transform method is the preferred choice because of its ability to avoid aliasing. It is commonly used in audio processing, communications, and other applications where preserving the filter’s frequency response is critical [155]. The impulse invariant method is useful in scenarios where matching the time-domain response of the analog filter is more important than avoiding aliasing, such as in certain control systems [184].
@@ 5.4.4 Additional Example: High-Pass Filter Design
To further illustrate the design of IIR filters, let’s design a high-pass filter using the bilinear transform method. High-pass filters allow high-frequency signals to pass while attenuating low-frequency signals. The steps are similar to the low-pass filter design but with a different filter type [136].
@@
14 plt.xlabel(’Normalized Frequency (×π rad/sample)’)
@@ 5.5 Stability and Causality in Filters
Filters are essential components in digital signal processing. They allow us to manipulate signals by emphasizing or attenuating certain frequency components. When designing filters, two critical properties must be considered: stability and causality. These properties ensure that the filter behaves correctly in real-world applications, maintaining a stable and predictable output [136].
@@ 5.5.1 Poles and Zeros
In the analysis of digital filters, the frequency response and stability of the system are largely determined by the location of its poles and zeros in the Z-plane [184]. The Z-plane is a complex plane where the poles and zeros of the system’s transfer function are plotted. The transfer function of a discrete-time system is typically expressed as a ratio of two polynomials [155]:
@@
  • The poles are the values of $z$ that make the denominator $A(z)$ equal to zero. The poles determine the frequencies where the filter’s response becomes unbounded or amplified [136].
The placement of poles and zeros in the Z-plane directly influences the filter’s frequency response [184]:
  • If the poles are close to the unit circle ($|z|=1$), the filter will have a more pronounced response at certain frequencies [155].
  • Zeros close to the unit circle result in attenuation at those frequencies [155].
Stability Condition: For the filter to be stable, all of the poles must lie strictly inside the unit circle in the Z-plane, i.e., $|z|<1$. If any pole lies on or outside the unit circle, the filter will be unstable, which means the output will grow without bound for certain inputs [136].
@@
We can visualize the pole-zero plot using Python and PyTorch to demonstrate the stability of this filter [144].
@@ 5.5.2 Causality and Stability Conditions
Causality is a fundamental property in filter design. A causal system is one where the output at any given time depends only on the current and past input values, not on future values [155]. This means that for a system to be physically realizable, it must be causal.
@@
  • A system is causal if the impulse response $h[n]$ of the system is zero for all $n<0$. That is, the output at any time depends only on inputs from the current and past, not from the future [184].
  • In terms of the Z-transform, a system is causal if the region of convergence (ROC) of the transfer function $H(z)$ includes the unit circle and is typically exterior to the outermost pole [136].
Stability: A system is considered stable if its output remains bounded for any bounded input. Mathematically, this condition is known as bounded input, bounded output (BIBO) stability [155]. In terms of the Z-transform, a system is stable if all of the poles of the system’s transfer function lie inside the unit circle, as discussed in the previous section [136].
@@ 6.1 Decimation and Interpolation
Multirate signal processing involves the manipulation of signals at various sampling rates. This is especially useful when signals need to be processed at different resolutions or bandwidths. Two fundamental operations in multirate signal processing are decimation and interpolation [65].
@@ 6.1.1 Decimation
  • $M$ is the decimation factor (i.e., the ratio by which the sampling rate is reduced) [39].
Decimation is often implemented in two steps [75]:
@@ 6.1.2 Interpolation
Interpolation is the process of increasing the sampling rate of a signal. This is done by introducing new samples between the existing samples. Interpolation is commonly used in applications where higher resolution or a higher sampling rate is needed [39].
The interpolation process typically involves two steps [198]:
@@
    Upsampling: The original signal is upsampled by inserting $L-1$ zeros between every sample, where $L$ is the interpolation factor.
@@
    Low-pass Filtering: After upsampling, the signal is passed through a low-pass filter to smooth the signal and remove spectral replicas introduced during upsampling. The cutoff frequency of the low-pass filter should be $\frac{f_{s}}{2L}$ [75].
@@ 6.2 Multirate System Structures
In multirate signal processing, we often need specialized structures to efficiently implement decimation and interpolation. Some common structures include polyphase filter implementations, decimators, and interpolators. These structures allow for computationally efficient processing, especially in systems where different sampling rates are used [198].
@@ 6.2.1 Polyphase Filter Implementations
Polyphase filters are efficient structures used in multirate systems to perform filtering and sample rate conversion simultaneously. The polyphase decomposition of a filter splits the filter into multiple sub-filters, each operating on different phases of the input signal. This allows for reduced computational complexity by avoiding redundant computations [65].
@@ 6.2.2 Efficient Decimator and Interpolator Designs
  • Audio and video signal processing [39].
  • Communications systems where different bandwidths and data rates are used [198].
  • Radar and sonar systems where signals are sampled at different resolutions [163].
@@ 6.2.3 Computational Efficiency in Multirate Systems
The use of polyphase filters and efficient decimators/interpolators significantly reduces computational cost in multirate systems. This is especially important when processing large datasets or real-time signals, where reducing the number of filter operations can lead to significant performance improvements [75].
@@ 6.3.1 Subband Coding
In subband coding, the signal is decomposed into multiple frequency bands using filters, and each band is processed at a different sampling rate. This technique is often used in speech and audio compression, where the signal can be represented more efficiently by encoding different frequency bands separately [200].
@@ 6.3.2 Speech and Audio Coding
Multirate techniques are critical in speech and audio coding algorithms such as MP3, AAC, and speech codecs. These algorithms reduce the amount of data needed to represent a signal by processing the signal at different rates for different frequency bands. This allows for high-quality sound reproduction at lower bit rates [22].
@@ 6.3.3 Adaptive Filtering
In adaptive filtering, multirate processing is used to optimize the performance of filters that need to adapt in real-time to changing signal conditions. For example, in communications systems, adaptive filters are used to mitigate interference and improve signal quality [60].
@@ 6.3.4 Communications and Wireless Systems
In modern communication systems, signals are often transmitted and received at different sampling rates. Multirate techniques such as decimation, interpolation, and polyphase filtering are essential for bandwidth reduction, data compression, and efficient signal processing in wireless communications [57].
@@ 6.3.5 Image Processing
In image processing, multirate techniques are used for tasks such as image scaling, subband decomposition in wavelet-based compression (e.g., JPEG2000), and multiresolution analysis. Decimation and interpolation allow for efficient manipulation of images at different resolutions [58].
@@ 7.1.1 Definition and Types of Images
An image is a two-dimensional signal where the intensity or color at each point represents visual information [58]. Unlike one-dimensional signals such as audio, images have both spatial dimensions width and height which allow them to represent detailed structures like shapes, objects, and scenes. In digital image processing, we deal with discrete images, where the continuous nature of real-world scenes is approximated by a grid of individual points called pixels.
@@
A grayscale image consists of pixels where each pixel value represents a single intensity level, typically ranging from black to white. Grayscale images contain only luminance information, without color, making them ideal for tasks that do not require color detail, such as medical imaging and object detection [86]. The pixel values in grayscale images usually range from 0 (black) to 255 (white) for 8-bit images, giving 256 different intensity levels.
@@
Color images contain more information than grayscale images by representing colors using three primary components: red, green, and blue (RGB) [58]. Each pixel in an RGB image is composed of three values corresponding to the intensities of red, green, and blue. By combining these components, we can represent a wide spectrum of colors. Each channel is typically 8 bits, so an RGB image with 24-bit depth can represent $216256^{3}=16,777,$ 216 possible colors. Color images are commonly used in photography, video, and multimedia applications [47].
@@
A binary image is the simplest form of an image, where each pixel can take only one of two possible values, typically 0 (black) or 1 (white). Binary images are useful in applications like text recognition and object segmentation [108], where the goal is to distinguish between two distinct regions or classes, such as foreground and background.
@@
Multispectral images capture data in multiple bands of the electromagnetic spectrum beyond the visible range [165]. Each pixel in a multispectral image contains several values, each representing a different spectral band. These images are widely used in remote sensing, environmental monitoring, and satellite imagery [120]. For example, multispectral satellite images may include bands for visible light, infrared, and ultraviolet wavelengths to provide a comprehensive view of the Earth’s surface. Hyperspectral images extend this idea further by capturing hundreds of bands, providing even more detailed spectral information.

### v2

@@ 5.2.2 Frequency Sampling Method
The Frequency Sampling Method designs FIR filters by specifying the desired frequency response at a set of equally spaced frequency points and then computing the corresponding time-domain coefficients using the inverse DFT [proakis1996digital].
@@
    Select the Number of Frequency Samples: Choose $N$, the length of the filter, which determines the number of frequency samples [oppenheim2010dsp].
@@
    Specify the Desired Frequency Response: Define the desired frequency response $H_{d}[k]$ at the $N$ frequency points [oppenheim2010dsp].
@@
    Compute the Time-Domain Coefficients: Obtain the filter coefficients $h[n]$ by computing the inverse DFT of $H_{d}[k]$ [proakis1996digital]:
@@
    Implement the Filter: Use the calculated coefficients $h[n]$ in the FIR filter implementation [proakis1996digital].
@@ 5.4 IIR Filter Design Techniques
Infinite Impulse Response (IIR) filters are widely used in digital signal processing because of their efficiency and ability to provide sharp frequency responses with a relatively low filter order [smith2007introduction]. Unlike FIR filters, IIR filters have feedback, which gives them an infinite impulse response. In this section, we will discuss two fundamental techniques for designing IIR filters: the bilinear transform method and the impulse invariant method [oppenheim2010dsp].
@@ 5.4.1 Bilinear Transform Method
The bilinear transform is one of the most popular methods for designing IIR filters. It allows us to take an analog filter design (such as a Butterworth, Chebyshev, or Elliptic filter) and map it into the digital domain [proakis1996digital]. This method works by mapping the analog $s$-plane to the digital $z$-plane using the following transformation:
@@
The bilinear transform preserves the frequency response characteristics of the analog filter while avoiding aliasing (a problem where high frequencies in the analog system would otherwise appear as lower frequencies in the digital system) [smith2007introduction]. However, one drawback is frequency warping, where the relationship between analog and digital frequencies becomes nonlinear. This can be corrected by pre-warping the critical frequencies before applying the transformation [oppenheim2010dsp].
@@
18 plt.xlabel(’Normalized Frequency (Œπ rad/sample)’)
@@ 5.4.2 Impulse Invariant Method
The impulse invariant method is another technique for converting analog filters into digital filters [proakis1996digital]. The key idea behind this method is to match the impulse response of the analog filter to the impulse response of the digital filter. This technique is particularly effective for designing low-pass filters where it is important to retain the shape of the analog impulse response in the digital domain [smith2007introduction].
In the impulse invariant method, we first compute the impulse response of the analog filter by taking the inverse Laplace transform of its transfer function. We then sample this continuous-time impulse response to obtain the discrete-time impulse response, which is subsequently used to determine the digital filter’s transfer function [oppenheim2010dsp].
One major drawback of the impulse invariant method is that it can introduce aliasing because high-frequency components of the analog filter may be folded into lower frequencies in the digital domain [smith2007introduction]. Therefore, this method is best suited for low-pass filters where aliasing is less of a concern [proakis1996digital].
@@
17 plt.xlabel(’Normalized Frequency (Œπ rad/sample)’)
@@ 5.4.3 Comparison of Bilinear Transform and Impulse Invariant Methods
Both the bilinear transform and impulse invariant methods have their advantages and drawbacks. Understanding their differences will help determine when to use each method [oppenheim2010dsp].
  • Bilinear Transform: The bilinear transform avoids aliasing by warping the frequency axis. However, it introduces frequency warping, which must be compensated for, especially in high-frequency regions [proakis1996digital].
  • Impulse Invariant Method: The impulse invariant method preserves the shape of the analog filter’s impulse response, but it can introduce aliasing. This method is typically preferred for low-pass filters where aliasing is less of a concern [smith2007introduction].
@@
In most practical applications, the bilinear transform method is the preferred choice because of its ability to avoid aliasing. It is commonly used in audio processing, communications, and other applications where preserving the filter’s frequency response is critical [proakis1996digital]. The impulse invariant method is useful in scenarios where matching the time-domain response of the analog filter is more important than avoiding aliasing, such as in certain control systems [smith2007introduction].
@@ 5.4.4 Additional Example: High-Pass Filter Design
To further illustrate the design of IIR filters, let’s design a high-pass filter using the bilinear transform method. High-pass filters allow high-frequency signals to pass while attenuating low-frequency signals. The steps are similar to the low-pass filter design but with a different filter type [oppenheim2010dsp].
@@
14 plt.xlabel(’Normalized Frequency (Œπ rad/sample)’)
@@ 5.5 Stability and Causality in Filters
Filters are essential components in digital signal processing. They allow us to manipulate signals by emphasizing or attenuating certain frequency components. When designing filters, two critical properties must be considered: stability and causality. These properties ensure that the filter behaves correctly in real-world applications, maintaining a stable and predictable output [oppenheim2010dsp].
@@ 5.5.1 Poles and Zeros
In the analysis of digital filters, the frequency response and stability of the system are largely determined by the location of its poles and zeros in the Z-plane [smith2007introduction]. The Z-plane is a complex plane where the poles and zeros of the system’s transfer function are plotted. The transfer function of a discrete-time system is typically expressed as a ratio of two polynomials [proakis1996digital]:
@@
  • The poles are the values of $z$ that make the denominator $A(z)$ equal to zero. The poles determine the frequencies where the filter’s response becomes unbounded or amplified [oppenheim2010dsp].
The placement of poles and zeros in the Z-plane directly influences the filter’s frequency response [smith2007introduction]:
  • If the poles are close to the unit circle ($|z|=1$), the filter will have a more pronounced response at certain frequencies [proakis1996digital].
  • Zeros close to the unit circle result in attenuation at those frequencies [proakis1996digital].
Stability Condition: For the filter to be stable, all of the poles must lie strictly inside the unit circle in the Z-plane, i.e., $|z|<1$. If any pole lies on or outside the unit circle, the filter will be unstable, which means the output will grow without bound for certain inputs [oppenheim2010dsp].
@@
We can visualize the pole-zero plot using Python and PyTorch to demonstrate the stability of this filter [paszke2019pytorch].
@@ 5.5.2 Causality and Stability Conditions
Causality is a fundamental property in filter design. A causal system is one where the output at any given time depends only on the current and past input values, not on future values [proakis1996digital]. This means that for a system to be physically realizable, it must be causal.
@@
  • A system is causal if the impulse response $h[n]$ of the system is zero for all $n<0$. That is, the output at any time depends only on inputs from the current and past, not from the future [smith2007introduction].
  • In terms of the Z-transform, a system is causal if the region of convergence (ROC) of the transfer function $H(z)$ includes the unit circle and is typically exterior to the outermost pole [oppenheim2010dsp].
Stability: A system is considered stable if its output remains bounded for any bounded input. Mathematically, this condition is known as bounded input, bounded output (BIBO) stability [proakis1996digital]. In terms of the Z-transform, a system is stable if all of the poles of the system’s transfer function lie inside the unit circle, as discussed in the previous section [oppenheim2010dsp].
@@ 6.1 Decimation and Interpolation
Multirate signal processing involves the manipulation of signals at various sampling rates. This is especially useful when signals need to be processed at different resolutions or bandwidths. Two fundamental operations in multirate signal processing are decimation and interpolation [harris2022multirate].
@@ 6.1.1 Decimation
  • $M$ is the decimation factor (i.e., the ratio by which the sampling rate is reduced) [crochiere1981interpolation].
Decimation is often implemented in two steps [hogenauer1981economical]:
@@ 6.1.2 Interpolation
Interpolation is the process of increasing the sampling rate of a signal. This is done by introducing new samples between the existing samples. Interpolation is commonly used in applications where higher resolution or a higher sampling rate is needed [crochiere1981interpolation].
The interpolation process typically involves two steps [vaidyanathan1990multirate]:
@@
    Upsampling: The original signal is upsampled by inserting L1L-1 zeros between every sample, where $L$ is the interpolation factor.
@@
    Low-pass Filtering: After upsampling, the signal is passed through a low-pass filter to smooth the signal and remove spectral replicas introduced during upsampling. The cutoff frequency of the low-pass filter should be $\frac{f_{s}}{2L}$ [hogenauer1981economical].
@@ 6.2 Multirate System Structures
In multirate signal processing, we often need specialized structures to efficiently implement decimation and interpolation. Some common structures include polyphase filter implementations, decimators, and interpolators. These structures allow for computationally efficient processing, especially in systems where different sampling rates are used [vaidyanathan1990multirate].
@@ 6.2.1 Polyphase Filter Implementations
Polyphase filters are efficient structures used in multirate systems to perform filtering and sample rate conversion simultaneously. The polyphase decomposition of a filter splits the filter into multiple sub-filters, each operating on different phases of the input signal. This allows for reduced computational complexity by avoiding redundant computations [harris2022multirate].
@@ 6.2.2 Efficient Decimator and Interpolator Designs
  • Audio and video signal processing [crochiere1981interpolation].
  • Communications systems where different bandwidths and data rates are used [vaidyanathan1990multirate].
  • Radar and sonar systems where signals are sampled at different resolutions [richards2014fundamentals].
@@ 6.2.3 Computational Efficiency in Multirate Systems
The use of polyphase filters and efficient decimators/interpolators significantly reduces computational cost in multirate systems. This is especially important when processing large datasets or real-time signals, where reducing the number of filter operations can lead to significant performance improvements [hogenauer1981economical].
@@ 6.3.1 Subband Coding
In subband coding, the signal is decomposed into multiple frequency bands using filters, and each band is processed at a different sampling rate. This technique is often used in speech and audio compression, where the signal can be represented more efficiently by encoding different frequency bands separately [vetterli1995wavelets].
@@ 6.3.2 Speech and Audio Coding
Multirate techniques are critical in speech and audio coding algorithms such as MP3, AAC, and speech codecs. These algorithms reduce the amount of data needed to represent a signal by processing the signal at different rates for different frequency bands. This allows for high-quality sound reproduction at lower bit rates [brandenburg1999mp3].
@@ 6.3.3 Adaptive Filtering
In adaptive filtering, multirate processing is used to optimize the performance of filters that need to adapt in real-time to changing signal conditions. For example, in communications systems, adaptive filters are used to mitigate interference and improve signal quality [goodwin2014adaptive].
@@ 6.3.4 Communications and Wireless Systems
In modern communication systems, signals are often transmitted and received at different sampling rates. Multirate techniques such as decimation, interpolation, and polyphase filtering are essential for bandwidth reduction, data compression, and efficient signal processing in wireless communications [goldsmith2005wireless].
@@ 6.3.5 Image Processing
In image processing, multirate techniques are used for tasks such as image scaling, subband decomposition in wavelet-based compression (e.g., JPEG2000), and multiresolution analysis. Decimation and interpolation allow for efficient manipulation of images at different resolutions [gonzalez2002digital].
@@ 7.1.1 Definition and Types of Images
An image is a two-dimensional signal where the intensity or color at each point represents visual information [gonzalez2002digital]. Unlike one-dimensional signals such as audio, images have both spatial dimensions width and height which allow them to represent detailed structures like shapes, objects, and scenes. In digital image processing, we deal with discrete images, where the continuous nature of real-world scenes is approximated by a grid of individual points called pixels.
@@
A grayscale image consists of pixels where each pixel value represents a single intensity level, typically ranging from black to white. Grayscale images contain only luminance information, without color, making them ideal for tasks that do not require color detail, such as medical imaging and object detection [kimpe2007increasing]. The pixel values in grayscale images usually range from 0 (black) to 255 (white) for 8-bit images, giving 256 different intensity levels.
@@
Color images contain more information than grayscale images by representing colors using three primary components: red, green, and blue (RGB) [gonzalez2002digital]. Each pixel in an RGB image is composed of three values corresponding to the intensities of red, green, and blue. By combining these components, we can represent a wide spectrum of colors. Each channel is typically 8 bits, so an RGB image with 24-bit depth can represent $256^{3}=16,777,216$ possible colors. Color images are commonly used in photography, video, and multimedia applications [duan2020video].
@@
A binary image is the simplest form of an image, where each pixel can take only one of two possible values, typically 0 (black) or 1 (white). Binary images are useful in applications like text recognition and object segmentation [lienhart2002localizing], where the goal is to distinguish between two distinct regions or classes, such as foreground and background.
@@
Multispectral images capture data in multiple bands of the electromagnetic spectrum beyond the visible range [rocchini2007using]. Each pixel in a multispectral image contains several values, each representing a different spectral band. These images are widely used in remote sensing, environmental monitoring, and satellite imagery [ma2023multispectral]. For example, multispectral satellite images may include bands for visible light, infrared, and ultraviolet wavelengths to provide a comprehensive view of the Earth’s surface. Hyperspectral images extend this idea further by capturing hundreds of bands, providing even more detailed spectral information.